Force of attraction btw two point charges Q and -q separated by d metre is F when these charges are placed on two identical spheres of radius R=0.3d whoese centre are d metre apart the force of attraction btw them is

Grade:12

1 Answers

Rasika
5 years ago
FFe=k*Q*-Q/d^2 F=k*Q*-Q/1.6d^2 F=Fe/1.6 Less than Fe therefore the force is less than the force on the two point charges 
R=0.3d
Distance between two sphere is 0.4d